What is the largest possible value of the expression n/3 - 3/n where n is an integer from 2 to 10 inclusive? Show work.
Answer by stanbon(75887)

What is the largest possible value of the expression n/3 - 3/n where n is an integer from 2 to 10 inclusive? Show work.
----
Large number in a numerator makes the fraction large.
---
Large number in a denominator makes the fraction small.
----
Ans: 10/3 - 3/10 = (100-9)/30 = 91/30
===================
Cheers,
Stan H.
===================